A force of 3kg wt is just sufficient to pull a block of mass 4kg over horizontal surface.what is the angle of friction.

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Let the angle of friction = theta
Tan  theta = f/R = 3g/ 4g = ¾
 
Theta = tan^-1 (3/4)
 
Hope it helps
 
Regards
Arun